# TaskEval: Assessing Difficulty of Code Generation Tasks for Large Language Models

### v2

@@ 6. Discussion
Echoing previous studies, our paper highlights that multiple prompts is a necessity to properly assess and evaluate CodeLLMs (Anwar et al., 2024). Indeed, variation in the prompts can drastically alter the output as we saw, for instance, in the motivating examples in Section 2. In our case, it is even more important as we aimed to evaluate concepts at the task-level, i.e., as independent of the formulation of the prompt as possible. Similarly, we showed that difficulty as assessed by humans might not be a valid proxy for CodeLLMs. Thus, using difficulty as reported in code-games benchmarks, such as CodeContest (Li et al., 2022) might not apply when evaluating CodeLLMs. Beyond evaluation, from a user perspective, extra care needs to be taken when prompting a CodeLLM, especially if the output of a single generation will be used. Using prompt tuning methods can help eliminate some biases (Fernando et al., 2023), but it may not completely address them. As such, a trade-off should be reached between using the output of a single prompt and using too many to be leveraged effectively by a user. In that regard, the use of differential testing or partial oracles (Wang and Zhu, 2024) could be a promising venue of improvement.
@@
In our study, analyzing tasks through the prism of IRT allows for some insights into the benchmark under study. Identifying tasks with varying difficulty helps reveal potential shortcomings of CodeLLMs, especially when this concerns a whole topic as we saw in RQ2. From there, further analysis can be conducted to understand why a difficulty is present. For instance, we saw in Section 2, from Listing 1, the impact that (the lack of) example could have on CodeLLMs generating a correct output in this SQL task. This could potentially be automated to provide a streamlined process of identifying shortcomings or hard topics. Then, the shortcomings should be addressed, potentially by fine-tuning the CodeLLMs on similar tasks/certain prompt templates or exploiting in-context learning (such as few-shot learning). To further train CodeLLMs, one can generate new tasks using the extracted information with the help of an approach such as Self-Instruct (Wang et al., 2022a) which allows generating new tasks based on a subset of sampled tasks through a generator LLM. This subset of tasks can be, for instance, the identified topics, as the prompt similarity between tasks can ease the generation. However, only giving a list of tasks marked as easy/hard is likely not to allow us to generate new hard/easy tasks accurately. Indeed, the limited number of tasks as well as the lack of explanation on why a given task is hard/easy would mean LLM tasked with generating new tasks will tend to struggle to identify what makes a task hard/easy just from the prompt. Instead, using a combinatory approach where one asks the generator LLM to augment a hard task by adding a constraint, or by adding a constraint from a hard task to an easy one (i.e., some form of fuzzing), could provide better results. This would ultimately reduce the diversity of generated tasks, but increase the likelihood that the generated tasks are hard.
@@
On the other hand, discriminant can be used to further select tasks to distinguish between CodeLLMs. Indeed, if the goal is, for a given benchmark, to differentiate between CodeLLMs, only the most discriminant tasks are required, as these tasks are the ones for which there are the most differences between CodeLLMs. On the contrary, focusing on low discriminant tasks can also highlight shortcomings but of a different nature. In that case, different CodeLLMs will perform similarly to each other. If all CodeLLMs also struggle on these tasks, it can emphasize particular tasks for which CodeLLMs under study are not adequately aligned. For instance, in RQ2, from Figure 5, we saw that the tasks in the SQL request topic of ClassEval had an average difficulty of $0.44$ and an average discriminant of $0.61$. Since all models struggle with tasks on this topic, regardless of their capability, this underscores that SQL-based tasks can still pose significant challenges for CodeLLMs.
@@ 7. Related Works
Evaluating CodeLLMs: Evaluation of CodeLLMs in the most simple form involves using a benchmark and having a CodeLLMs under test generates code for each prompt in the benchmark. Then, an aggregated metric can be calculated to evaluate the performance of the CodeLLMs on the benchmark. Several benchmarks for code have been proposed over the years, each with some particularity: besides HumanEval+ and ClassEval used in this work, APPS (Hendrycks et al., 2021) is a benchmark based on coding competition code fragments, CodeXGlue (Lu et al., 2021) contains additional tasks beyond code generation such as code completion or code translation or BigCodeBench (Big, 2024b) which contains more complex instructions. The difference between those benchmarks mainly stem from the different levels of dependency needed to solve the task (e.g., simple function for HumanEval+ or class level for ClassEval), yet the evaluation mainly relies on aggregation over the whole benchmark using accuracy or pass@k. More recently, multiple benchmarks are being proposed to measure a particular property of the CodeLLMs by collecting/crafting in a particular way: Coignion et al. (Coignion et al., 2024) evaluated the speed of generated code based on LeetCode, Liu et al. (Liu et al., 2024b) created HalluCode to evaluate CodeLLMs attitude towards different types of hallucination or Jain et al. (Jain et al., 2024b) introducing LiveCode bench to measure contamination in CodeLLMs. However, the properties evaluated are generally at the benchmark level and the possible impact of the rephrasing is not evaluated. Moreover, when it comes to difficulty, most benchmarks make use of human judgment, for instance in the case of APPS, by using the subjective score given by code game contestants. We showed in RQ4 that this assessment might not translate to the actual difficulty of realizing the task for CodeLLMs. This result echoes to similar observations in other studies: Ouyang et al. (Ouyang et al., 2024) showed that the test pass rate does not necessarily correlate with difficulty, as aggregated by human participants in a coding contest, and Kou et al. (Kou et al., 2023) highlighted major differences when comparing how a human developer and CodeLLMs process codes’ instruction.
@@
Item Response Theory to evaluate NLP models: IRT (Baker, 2001) has been used extensively to evaluate NLP models. Sedoc et al. (Sedoc and Ungar, 2020) used IRT to compare chatbots’ proficiency by leveraging human annotators’ judgment on the chatbots’ output as observations. They notably show that the IRT model can be used to further streamline evaluations by selecting the most discriminative examples. Rodriguez et al. (Rodriguez et al., 2021b) propose to leverage IRT to redefine leaderboards in NLP model evaluation by defining the Difficulty and Ability Discriminating (DAD) approach. They show that IRT models offer greater reliability in assessing the ranking stability of questions both within and across benchmarks and can provide valuable qualitative insights into the evaluated leaderboards. Similarly, Vania et al. (Vania et al., 2021b) uses a unique IRT model to analyze 29 NLP datasets and 18 Transformers, producing insights into which dataset contains more difficult examples or which is more effective at distinguishing strong models. Zhuang et al. (Zhuang et al., 2024) explain how IRT can be used for adaptive testing of CodeLLMs highlighting the limitations of traditional metrics such as Accuracy. First, compared to those studies, we focus on coding tasks and, more particularly, code generation. Secondly, these studies consider an analysis of the benchmark prompts. On the contrary, we aim at measuring the difficulty of the task itself, as previous studies showed that prompt formulation has a drastic impact on the models’ output (Mizrahi et al., 2024; Weber et al., 2023; Wang et al., 2022b).
@@ 8. Threats to Validity
Construct validity: To derive the characteristics of the tasks we rely on IRT which has been extensively used in education measurement (Zanon et al., 2016), as well as in assessing LLM in classical NLP settings (Rodriguez et al., 2021a; Lalor et al., 2016). As such, the backbone of our approach is grounded in theory. We used a multi-prompted approach as previous studies highlighted the importance of multiple phrasing. To avoid potential local dependence effect (i.e., measure a score using multiple prompts referring to the same tasks), we aggregate results of prompts for a same task, effectively building polytomous super-items or “testlets” (Wainer and Kiely, 1987). We rely on three levels of prompts to model the effect of information on how the CodeLLMs handle the task. The choice of three levels was made based on our observations. So, while we limited our study to three levels of prompts, we do not expect providing more information in prompts affects our results drastically. Another threat might come from the generation of the prompts. The transformations used are motivated based on existing works. Moreover, while we have made use of state-of-the-art GPT-4 to generate them, we manually checked the prompts to make sure that the generative process did not alter semantics. Finally, another threat to validity comes from the fact that we did not consider examples in the prompts. This, as we mentioned, can decrease the performance of the models and so add artificially hard tasks. This was nonetheless necessary to make the comparison fair across tasks as not all examples contain the same information, as we gave some examples in Section 2. However, beyond modifying some tasks’ characteristics, it does not affect the applicability of TaskEval.
@@
Internal validity: The CodeLLMs and benchmarks selected could be a threat to validity. We made sure to select different architectures through $5$ different CodeLLMs to encompass different models. We used the implementation and models’ weights as present on HuggingFace (HuggingFace, [n. d.]) by the original authors of each model or the web API in the case of GPT3.5. For the benchmarks, we selected two different datasets, both acting at different levels (functional for HumanEval+ and class for ClassEval). Regarding our experiments, the modeled topics as well as the human-annotations data could be a limitation. For the topics, we used BerTopics to automatically label the tasks, with noisy tasks being discarded. This could reduce our potential topics list, but this allowed for an automatic process. We still obtained over 15 topics with at least $3$ tasks each. We manually reviewed the obtained topics. Regarding human annotations, the data primarily reflect how difficult annotators perceive each task to be. As detailed in RQ4, we did not ask crowd-workers to code the tasks, as this could introduce biases due to the uncontrolled use of CodeLLMs. Instead we followed an existing approach (Alhamed and Storer, 2021) based on effort estimation which mitigates this issue. Although the difficulty is then assessed based on developers’ subjective judgment, the used approach allows to obtain an estimation of a consensus which limits potential outliers due to different experience, relative estimation etc. Moreover, even under variations due to human judgment, participants’ experience etc. we still show there is a delta between human’s and LLM’s perception of the difficulty, which further motivate proposing an approach that is LLM-centric.
@@
External validity: Our selection of tasks from the chosen benchmarks may pose a threat to validity, as the collection might not fully represent real-world programming tasks for assessing LLMs’ capabilities. To mitigate this, we used two benchmarks, each focusing on different aspects. Moreover, both HumanEval+ and ClassEval have been used to evaluate CodeLLM performance (Eva, 2024; Cla, 2024) in previous studies. Finally, this study aimed to establish a foundation for assessing task characteristics rather than focusing on their practical implementation. We believe that using tasks from these two benchmarks offers a diverse set of use cases for evaluating code generation in CodeLLMs. The approach should not be affected by more practical benchmarks, and we leave to future works the application of the method to said benchmarks.
@@ 9. Conclusion
In this paper, we present a method for analyzing the characteristics of code generation tasks for CodeLLMs. Specifically, we introduce TaskEval, a novel framework designed to assess and explore task characteristics in depth. We evaluated TaskEval on two well-known benchmarks for CodeLLM-based code generation: HumanEval+ and ClassEval. A set of diverse prompts is crafted for each task including different levels of contextual information about the task and various rephrasings, accounting for the impact of prompts variations on the generation. Then, five state-of-the-art CodeLLMs were used to solve coding tasks using the set of prompts. TaskEval is based on Item Response Theory and allows to compute difficulty (i.e., how hard is a given task for CodeLLMs) and discriminant (i.e., how do CodeLLMs compare to each other on a given task) of each task within a benchmark. Using our method, we conducted multiple analyses on the tasks, namely topic-wise, program construct-wise, as well as a comparison with human annotators. Based on thematic similarities, we clustered tasks from both benchmarks into respectively 17 and 21 topics to inspect their characteristics, revealing topics with high difficulty and high discriminant for CodeLLMs (e.g., Sequences). Similarly, we assess program construct (e.g., conditions, variable assignment etc.) used given the tasks’ difficulty/discriminant and highlighted discrepancy/similarity between benchmarks. Finally, we compared the difficulty assessment between CodeLLMs and human annotators, revealing stark differences in the assessment. In future work, we plan to extend the evaluation of TaskEval to additional domains, such as code completion and summarization. Furthermore, we aim to build on our analysis to enhance task selection (e.g., based on discriminative features) and facilitate task generation (e.g., using topic modeling).

### v3

@@ 6. Discussion
With TaskEval, we showed how to leverage IRT’s outcome to analyse benchmarks to provide a more fine-grained analysis. While the aim of our study was not to primarily focus on addressing practical aspects such as computational cost or scalability, we highlighted several key practical points and considerations that could benefit practitioners.
@@
Multi-prompt evaluation: Echoing previous studies, we re-emphasise that multiple prompts are a necessity to assess and evaluate CodeLLMs (Anwar et al., 2024) properly. Indeed, variation in the prompts can drastically alter the output as we saw, for instance, in the motivating examples in Section 2 with the Pass@1 drastically changing depending on the rephrasing and/or level of information in the prompt. In our case, this was even more important as we aimed to evaluate concepts at the task-level, i.e., as independent of the formulation of the prompt as possible. Beyond evaluation, from a user perspective, extra care needs to be taken when prompting a CodeLLM, especially if the output of a single generation will be used. Using prompt tuning methods can help eliminate some biases (Fernando et al., 2023), but it may not completely address them. As such, a trade-off should be reached between using the output of a single prompt and using too many to be leveraged effectively by a user. In that regard, the use of differential testing or partial oracles (Wang and Zhu, 2024) could be a promising venue of improvement.
Human assessment of difficulty: We showed that difficulty as assessed by humans might not be a valid proxy for CodeLLMs, and so a misalignment between human and CodeLLMs task difficulty might exist. Thus, using difficulty as reported in code-games benchmarks, such as CodeContest (Li et al., 2022), might not apply when evaluating CodeLLMs. Leveraging instead the difficulty as computed by an IRT model from the CodeLLMs’ outputs themselves might be more truthful to the real difficulty of the task for CodeLLMs. This observation also sheds light on a broader cautionary point: since benchmarks are often used to assess CodeLLM capabilities, potentially in relation to developers’ skills, if they mischaracterise what is simple or difficult for a CodeLLM, claims about CodeLLMs might be overstated.
@@
Identifying CodeLLMs shortcomings: In our study, analysing tasks through the prism of IRT allows for some insights into the benchmark under study. Identifying tasks with varying difficulty helps reveal potential shortcomings of CodeLLMs, especially when this concerns a whole topic, as we saw in RQ2. From there, further analysis can be conducted to understand why a difficulty is present. This could potentially be automated to provide a streamlined process of identifying shortcomings or hard topics. Then, the shortcomings should be addressed, potentially by fine-tuning the CodeLLMs on similar tasks/certain prompt templates or exploiting in-context learning (such as few-shot learning). To further train CodeLLMs, one can generate new tasks using the extracted information with the help of an approach such as Self-Instruct (Wang et al., 2022a), which allows generating new tasks based on a subset of sampled tasks through a generator LLM. This subset of tasks can be, for instance, the identified topics, as the prompt similarity between tasks can ease the generation. However, only giving a list of tasks marked as easy/hard is likely not to allow us to generate new hard/easy tasks accurately. Indeed, the limited number of tasks as well as the lack of explanation on why a given task is hard/easy would mean LLM tasked with generating new tasks will tend to struggle to identify what makes a task hard/easy just from the prompt. Instead, using a combinatory approach where one asks the generator LLM to augment a hard task by adding a constraint, or by adding a constraint from a hard task to an easy one (i.e., some form of fuzzing), could provide better results. This would ultimately reduce the diversity of generated tasks, but increase the likelihood that the generated tasks are hard.
@@
Evaluating benchmark discriminability: Discriminant can be used to further select tasks to distinguish between CodeLLMs. Indeed, if the goal is, for a given benchmark, to differentiate between CodeLLMs, only the most discriminant tasks are required, as these tasks are the ones for which there are the most differences between CodeLLMs. On the contrary, focusing on low discriminant tasks can also highlight shortcomings, but of a different nature. In that case, different CodeLLMs will perform similarly to each other. If all CodeLLMs also struggle on these tasks, it can emphasise particular tasks for which CodeLLMs under study are not adequately aligned. For instance, in RQ2, from Figure 5, we saw that the tasks in the Condition checking or Data structure calculation topics of ClassEval had an average difficulty of 0.4 and 0.64 and an average discriminant of 0.65 and 1. Since all models can somehow struggle with tasks on these topics, regardless of their capability, this underscores that such tasks can still pose significant challenges for CodeLLMs.
Cost and Scalability: In terms of cost, the most expensive part in our experiments was prompt generation and verification. The overall cost is proportional to the number of benchmarks, tasks, and transformations applied. For the prompt generation, using GPT-4 turbo, this amounted to  50\$ for all tasks and all levels ($\sim$6,500 prompts), accounting for rerun when correcting them, resulting in 0.007\$/prompt. Note that GPT-4 turbo is an older model and the API cost has since decreased drastically on newer versions, which could easily divide the cost of the experiments we did. For the prompt verification, this had to be done manually, which translated to several dozen hours of work. This was done to ensure as few errors as possible in the generated prompts, and is a limiting factor for the scalability that could be a future work opportunity. Note this, however, needs only to be done once for the benchmarks under study. For the generation, the cost is proportional to the number of models (8 in our case) and the number of inferences ($\sim$32,$0$ in our case). In our case, the open-source models used were run on a single 32GB VRAM GPU with less than an hour of runtime each. For the closed-source model (GPT3.5 and DeepSeek-V3), the cost was  5\$ each, for all tasks and all prompts, resulting in 0.0002\$/inference. IRT model training runtime/cost is negligible. Overall, we believe the approach is relatively inexpensive in terms of cost, with only the manual verification being a limiting factor, and so could be easily scalable, either through application to new benchmarks or to more models.
@@ 7. Related Works
Evaluating CodeLLMs: Evaluation of CodeLLMs in the most simple form involves using a benchmark, and having a CodeLLM under test generates code for each prompt in the benchmark. Then, an aggregated metric can be calculated to evaluate the performance of the CodeLLMs on the benchmark. Several benchmarks for code have been proposed over the years, each with some particularity: besides HumanEval+ and ClassEval used in this work, APPS (Hendrycks et al., 2021) is a benchmark based on coding competition code fragments, CodeXGlue (Lu et al., 2021) contains additional tasks beyond code generation, such as code completion or code translation or BigCodeBench (Big, 2024b), which contains more complex instructions. The difference between those benchmarks mainly stems from the different levels of dependency needed to solve the task (e.g., a simple function for HumanEval+ or class level for ClassEval), yet the evaluation mainly relies on aggregation over the whole benchmark using accuracy or pass@k. More recently, multiple benchmarks are being proposed to measure a particular property of the CodeLLMs by collecting/crafting in a particular way: Coignion et al. (Coignion et al., 2024) evaluated the speed of generated code based on LeetCode, Liu et al. (Liu et al., 2024b) created HalluCode to evaluate CodeLLMs’ attitude towards different types of hallucination, or Jain et al. (Jain et al., 2024) introduced LiveCode bench to measure contamination in CodeLLMs. However, the properties evaluated are generally at the benchmark level, and the possible impact of the rephrasing is not evaluated. Moreover, when it comes to difficulty, most benchmarks make use of human judgment, for instance, in the case of APPS, by using the subjective score given by code game contestants. We showed in RQ4 that this assessment might not translate to the actual difficulty of realising the task for CodeLLMs. This result echoes similar observations in other studies: Ouyang et al. (Ouyang et al., 2024) showed that the test pass rate does not necessarily correlate with difficulty, as aggregated by human participants in a coding contest, and Kou et al. (Kou et al., 2023) highlighted major differences when comparing how a human developer and CodeLLMs process code instructions.
@@
Item Response Theory to evaluate NLP models: IRT (Baker, 2001) has been used extensively to evaluate NLP models. Sedoc et al. (Sedoc and Ungar, 2020) used IRT to compare chatbots’ proficiency by leveraging human annotators’ judgment on the chatbots’ output as observations. They notably show that the IRT model can be used to further streamline evaluations by selecting the most discriminative examples. Rodriguez et al. (Rodriguez et al., 2021b) propose to leverage IRT to redefine leaderboards in NLP model evaluation by defining the Difficulty and Ability Discriminating (DAD) approach. They show that IRT models offer greater reliability in assessing the ranking stability of questions both within and across benchmarks and can provide valuable qualitative insights into the evaluated leaderboards. Similarly, Vania et al. (Vania et al., 2021b) use a unique IRT model to analyse 29 NLP datasets and 18 Transformers, producing insights into which dataset contains more difficult examples or which is more effective at distinguishing strong models. Zhuang et al. (Zhuang et al., 2024) explain how IRT can be used for adaptive testing of CodeLLMs, highlighting the limitations of traditional metrics such as Accuracy. First, compared to those studies, we focus on coding tasks and, more particularly, code generation. Secondly, these studies consider an analysis of the benchmark prompts. On the contrary, we aim at measuring the difficulty of the task itself, as previous studies showed that prompt formulation has a drastic impact on the models’ output (Mizrahi et al., 2024; Weber et al., 2023; Wang et al., 2022b). Moreover, our approach differs from prompt engineering techniques (Sasaki et al., 2024) as they aim to tune the model/prompt to achieve the best result on a particular task. In contrast, we generate several prompts to represent the diversity of possible CodeLLM answers for a particular task and observe the result.
@@ 8. Threats to Validity
Construct validity: To derive the characteristics of the tasks, we rely on IRT, which has been extensively used in education measurement (Zanon et al., 2016), as well as in assessing LLM in classical NLP settings (Rodriguez et al., 2021a; Lalor et al., 2016). IRT assumes local independence, which might be jeopardised in our case because we use multiple prompts for the same task. To avoid a potential local dependence effect, we aggregate results of prompts for the same task, effectively building polytomous superitems or “testlets” (Wainer and Kiely, 1987; Wainer et al., 2007). This was shown to alleviate the local dependence effect. As such, the backbone of our approach is grounded in theory. We used a multi-prompted approach as previous studies highlighted the importance of multiple phrasings. We rely on three levels of prompts to model the effect of information on how the CodeLLMs handle the task. The choice of three levels was made based on our observations. So, while we limited our study to three levels of prompts, we do not expect that providing more information in prompts will drastically affect our results. Another threat might come from the generation of the prompts. The transformations used are motivated based on existing works. Moreover, while we have made use of state-of-the-art GPT-4 to generate them, we manually checked the prompts to make sure that the generative process did not alter semantics. Choosing GPT4 for prompt generation might unintentionally bias the evaluation by benefiting a model trained similarly (GPT3.5). However, observing our results, especially the reported abilities by the trained IRT models, seems to point towards this not influencing the results. Indeed, the abilities obtained reflect the “perceived” capacity of the model. Finally, we did not consider examples in the prompts. This, as we mentioned, can decrease the performance of the models and so add artificially hard tasks. This was nonetheless necessary to make the comparison fair across tasks, as not all examples contain the same information, as we gave some examples in Section 2. However, beyond modifying some tasks’ characteristics, it does not affect the applicability of TaskEval.
@@
Internal validity: The CodeLLMs and benchmarks selected could be a threat to validity. We made sure to select different architectures through $8$ different CodeLLMs to encompass different models. We used the implementation and models’ weights as present on HuggingFace (HuggingFace, [n. d.]) by the original authors of each model or the web API in the case of GPT3.5 and DeepSeek-V3. For the benchmarks, we selected two different datasets, both acting at different levels (functional for HumanEval+ and class for ClassEval). Regarding our experiments, the modelled topics as well as the human-annotation data could be a limitation. For the topics, we used BERTopics to automatically label the tasks, with noisy tasks being discarded. This could reduce our potential topics list, but this allowed for an automatic process. We still obtained over 15 topics with at least $3$ tasks each. We manually reviewed the obtained topics. Regarding human annotations, the data primarily reflect how difficult annotators perceive each task to be. As detailed in RQ4, we did not ask crowd-workers to code the tasks, as this could introduce biases due to the uncontrolled use of CodeLLMs. Instead, we followed an existing approach (Alhamed and Storer, 2021) based on effort estimation, which mitigates this issue. Although the difficulty is then assessed based on developers’ subjective judgment, the used approach allows for obtaining an estimation of a consensus, which limits potential outliers due to different experiences, relative estimation, etc. Moreover, even under variations due to human judgment, participants’ experience, etc., we still show there is a delta between humans’ and CodeLLM’s perception of the difficulty, which further motivates proposing an approach that is LLM-centric.
@@
External validity: Our selection of tasks from the chosen benchmarks may pose a threat to validity, as the collection might not fully represent real-world programming tasks for assessing LLMs’ capabilities. To mitigate this, we used two benchmarks, HumanEval+ and ClassEval, which have been used to evaluate CodeLLM performance (Eva, 2024; Cla, 2024) in previous studies. Both datasets cover different functionality levels (function for HumanEval+ and class for ClassEval). We chose their Python version as it is the most widespread programming language when it comes to LLMs (Jiang et al., 2024). Finally, this study aimed to establish a foundation for assessing task characteristics, as well as showing that interesting insights can be gathered from the analysis. The approach itself should not be affected by more practical benchmarks. Nonetheless, this limits the generalisation of our study in terms of practical insights. We had to limit ourselves to these two datasets to reduce the complexity of both the analysis and the cost of the prompt generation/human judgement experiments. Studying more advanced benchmarks (e.g. SWE-bench (Jimenez et al., 2024)), benchmarks using other programming languages (e.g. Multi-Eval (Athiwaratkun et al., 2022)) or industry-oriented benchmarks would add valuable insights to the capacity of LLMs on a more diverse panel of tasks, and so is an important future venue of research.
@@ 9. Conclusion
In this paper, we present a method for analysing the characteristics of code generation tasks for CodeLLMs. Specifically, we introduce TaskEval, a novel framework designed to assess and explore task characteristics in depth. We evaluated TaskEval on two well-known benchmarks for CodeLLM-based code generation: HumanEval+ and ClassEval. A set of diverse prompts is crafted for each task, including different levels of contextual information about the task and various rephrasings, accounting for the impact of prompt variations on the generation. Then, eight CodeLLMs were used to solve coding tasks using the set of prompts. TaskEval is based on Item Response Theory and allows for computing difficulty (i.e., how hard is a given task for CodeLLMs) and discriminant (i.e., how do CodeLLMs compare to each other on a given task) of each task within a benchmark. Using our method, we conducted multiple analyses on the tasks, namely, topic-wise, program construct-wise, as well as a comparison with human annotators. Based on thematic similarities, we clustered tasks from both benchmarks into 17 and 21 topics to inspect their characteristics, revealing topics with high difficulty and high discriminant for CodeLLMs (e.g., Sequences). Similarly, we assess program construct (e.g., conditions, variable assignment, etc.) given the tasks’ difficulty/discriminant and highlighted discrepancies/similarities between benchmarks. Finally, we compared the difficulty assessment between CodeLLMs and human annotators, revealing stark differences in the assessment. In future work, we plan to extend the evaluation of TaskEval to additional domains, such as code completion and summarisation. Furthermore, we aim to build on our analysis to enhance task selection (e.g., based on discriminative features) and facilitate task generation (e.g., using topic modelling).
Acknowledgments
This work was partly supported by: Fonds de Recherche du Québec (FRQ), the Canadian Institute for Advanced Research (CIFAR) as well as the DEEL project CRDPJ 537462-18 funded by the Natural Sciences and Engineering Research Council of Canada (NSERC) and the Consortium for Research and Innovation in Aerospace in Québec (CRIAQ), together with its industrial partners Thales Canada inc, Bell Textron Canada Limited, CAE inc and Bombardier inc.